### v2

# MnLargeSymbols’164 MnLargeSymbols’171

Abstract
The choice of batch sizes in minibatch stochastic gradient optimizers is critical in large-scale model training for both optimization and generalization performance. Although large-batch training is arguably the dominant training paradigm for large-scale deep learning due to hardware advances, the generalization performance of the model deteriorates compared to small-batch training, leading to the so-called “generalization gap” phenomenon. To mitigate this, we investigate adaptive batch size strategies derived from adaptive sampling methods, originally developed only for stochastic gradient descent. Given the significant interplay between learning rates and batch sizes, and considering the prevalence of adaptive gradient methods in deep learning, we emphasize the need for adaptive batch size strategies in these contexts. We introduce AdAdaGrad and its scalar variant AdAdaGrad-Norm, which progressively increase batch sizes during training, while model updates are performed using AdaGrad and AdaGrad-Norm. We prove that AdAdaGrad-Norm converges with high probability at a rate of $\mathscr{O}(1/K)$ to find a first-order stationary point of smooth nonconvex functions within $K$ iterations. AdAdaGrad also demonstrates similar convergence properties when integrated with a novel coordinate-wise variant of our adaptive batch size strategies. We corroborate our theoretical claims by performing image classification experiments, highlighting the merits of the proposed schemes in terms of both training efficiency and model generalization. Our work unveils the potential of adaptive batch size strategies for adaptive gradient optimizers in large-scale model training.
1 Introduction
Large-scale optimization algorithms (Bottou et al., 2018) form the foundation of the success of deep learning in the era of generative AI. Minibatch stochastic gradient descent (SGD) (Robbins and Monro, 1951) and its many variants, along with batch sampling techniques, are the main workhorse for training deep neural networks. However, training deep neural networks, such as those in transformers, is notoriously challenging because of their high dimensionality and nonconvex landscape. This complexity necessitates extensive hyperparameter tuning and sophisticated training strategies for optimizers to circumvent premature divergence and training instabilities. Consequently, training deep learning models often appears as more an art than a science. The most critical hyperparameter is arguably the learning rate (or step size). Adaptive gradient methods with adaptive learning rates, such as AdaDelta (Zeiler, 2012), AdaGrad (Duchi et al., 2011), and Adam (Kingma and Ba, 2015), are now prevalent, as they reduce the need for meticulous tuning and complex learning rate schedules, typically required for SGD. Another important, yet frequently overlooked hyperparameter is the batch size. It moderates the trade-off between computational efficiency and model generalization by controlling the magnitude of noise in batch gradients. However, the selection of batch sizes in deep learning training remains largely heuristic, such as using a constant size for convolutional networks or a linear warm-up for large language models (Brown et al., 2020; Rae et al., 2021; Hoffmann et al., 2022), and is usually predetermined before training begins. Furthermore, from a hardware utilization perspective, the use of a large number of distributed computational resources (i.e., GPUs or TPUs) necessitates the large-batch training regime when parallel minibatch SGD (Zinkevich et al., 2010; Dean et al., 2012) is used. In addition, the intricate relationship between learning rates and batch sizes deserves attention. Specifically, Smith et al. (2018) demonstrated the equivalence of reducing step sizes and increasing batch sizes, but this principle applies mainly to SGD. The impact of varying batch sizes in adaptive gradient methods has not yet been fully explored.
In light of this, our work seeks to unravel the complexities surrounding the determination of optimal batch sizes for adaptive gradient methods. We aim to introduce automated schedules capable of dynamically deciding when to increase the batch size and by how much, based on the training needs. Our approach is theoretically grounded and relies on the statistics of batch gradients of iterates, thereby adapting to the training dynamics of the models. Our proposed schedules are able to reduce the generalization gap in large-batch training, but at the same time also maintain the utilization of large batches at the later stage of the overall training process. Our proposed strategies are based on adaptive sampling methods (Byrd et al., 2012; Bollapragada et al., 2018a). In the context of deep learning, De et al. (2016, 2017) have numerically shown the effectiveness of these methods with AdaDelta. However, the convergence properties of such adaptive sampling methods for adaptive gradient methods have not been thoroughly investigated, leaving a gap between theory and practice. Moreover, in existing adaptive sampling methods developed mainly for SGD, step sizes are often fixed or adjusted using backtracking line search procedures (mainly for convex problems). This becomes computationally impractical or inefficient for larger models, especially given the nonconvex nature of deep neural networks.
The development of adaptive batch size strategies for deep learning is not a novel concept, featuring methodologies such as Big Batch SGD (De et al., 2016, 2017), CABS (Balles et al., 2017), AdaBatch (Devarakonda et al., 2017), SimiGrad (Qin et al., 2021), and AdaScale SGD (Johnson et al., 2020), which adapts learning rates for large batches rather than directly adjusting batch sizes. However, these methods often lack a principled basis with rigorous convergence guarantees or are limited to analyses of SGD under stringent conditions (e.g., convexity or the Polyak–Łojasiewicz condition), despite the prevalence of adaptive gradient methods in deep learning. Furthermore, these approaches still necessitate the choice of learning rate adjustments, such as line search routines or schedulers, leaving a gap in achieving full adaptivity. Furthermore, strategies for determining new batch sizes may rely on heuristic rules, such as geometric growth or decay rates (Qin et al., 2021).
Contributions.
In this work, our objective is to bridge the theoretical and practice gaps of adaptive batch size schemes for adaptive gradient methods. We introduce two principal adaptive batch size schemes, grounded in the so-called adaptive sampling methods (Byrd et al., 2012; Friedlander and Schmidt, 2012; Bollapragada et al., 2018a), tailored for adaptive gradient methods. Our focus is on AdaGrad and its norm variant AdaGrad-Norm, which are among the simplest and most extensively studied adaptive gradient methods. The development of adaptive batch size schemes for these methods has significant theoretical and practical relevance.
The technical contributions of this work are three-fold. From a theoretical perspective, we establish a sublinear convergence rate (with high probability) for our proposed methods when applied with AdaGrad-Norm and AdaGrad to smooth nonconvex objectives. This significantly broadens the existing body of work which has mainly focused on SGD. Moreover, we relax the Lipschitz smoothness condition on the objective function by adopting the generalized smoothness concept introduced in Zhang et al. (2020b, a). This adaptation allows for more generalized and realistic applications in the current practice of deep learning. On the empirical front, we demonstrate the efficacy of our proposed methods through various numerical experiments on image classification tasks. These experiments highlight the benefits of the adaptivity of our schemes, driven by both adaptive batch sizes and step sizes. Finally, we provide an efficient implementation of the proposed approach in PyTorch, making use of the torch.func module for efficient parallelized computations of per-sample gradients using the so-called vectorizing map function vmap. An open source code repository will be provided upon acceptance. To the best of our knowledge, our proposed methods are the first adaptive batch size schemes based on adaptive sampling methods for adaptive gradient methods, which are convergent with guarantees, empirically performant and efficiently implemented in deep learning libraries.
2 Related Work
We give an overview of related work on batch sizes in model training and adaptive sampling methods. Further related work can be found in Appendix A.
Large-batch training.
In stochastic gradient optimizers, batch sizes play a crucial role in controlling the variance of stochastic gradients as estimators of the full deterministic gradients. Although the noise from stochastic gradients offers benefits for nonconvex optimization, the trend towards using larger batches in large-scale model training has become standard, thanks to advancements in parallelized hardware such as GPUs, significantly reducing training time. The concept of large-batch training in deep learning was popularized in Smith and Le (2018); Smith et al. (2018) and has been widely adopted in applications such as ImageNet classification (Goyal et al., 2017) and BERT training (You et al., 2020). In the work of LeCun et al. (2002), it has been understood that model generalization deteriorates in large-batch training. Large-batch training tends to result in a loss landscape with many sharp minima, which are harder to escape during optimization and hence lead to worse generalization (Keskar et al., 2017). The impact of batch sizes has been further examined in a more principled manner in subsequent studies. For example, McCandlish et al. (2018) introduced an empirical model for large batch training without providing rigorous proof, postulating the existence of critical batch sizes through extensive numerical simulations on convolutional neural networks (CNNs), LSTMs, and VAEs. Meanwhile, Kaplan et al. (2020) focused their experiments on transformers. Zhang et al. (2019) explored how critical batch sizes vary with the nature of optimizers, including acceleration (momentum), preconditioning, and exponential moving average, through both large-scale experiments and a noisy quadratic model. Granziol et al. (2022) applied random matrix theory to examine the batch Hessian, theoretically establishing learning rates as a function of batch size (i.e., the scaling rule) for both SGD (linear) and adaptive gradient methods (square root), with experimental validation. Although the “generalization gap” can be narrowed by using larger learning rates proportional to batch sizes to maintain the gradient noise scale (Hoffer et al., 2017; Smith and Le, 2018; Smith et al., 2018), it cannot be completely eliminated. However, Shallue et al. (2019) investigated the impact of batch sizes in the context of data parallelism and empirically characterized the effects of large-batch training, finding no evidence of a degradation in generalization performance.
Adaptive sampling methods.
In the realm of stochastic optimization, a more theoretically grounded approach known as adaptive sampling methods has been developed for batch (or minibatch) algorithms. Byrd et al. (2012) introduced a method termed the norm test, which adaptively increases the batch size throughout the optimization process. The rationale behind the norm test traces back to Carter (1991) and represents a more generalized condition compared to the approach of Friedlander and Schmidt (2012). Both approaches demonstrate linear convergence when the batch size increases geometrically. Bollapragada et al. (2018a) proposed the augmented inner product test, which allows for more gradual increases in batch sizes than the norm test. Furthermore, Cartis and Scheinberg (2018) introduced a relaxation of the norm test that allows its condition to be violated with a probability of less than $0.5$. This suite of adaptive sampling methods is considered part of the variance-reduced optimization algorithms (Johnson and Zhang, 2013; Reddi et al., 2016) that are widely used in machine learning (Gower et al., 2020).
3 Problem Formulation
Notation.
[…]
Problem setting.
We consider the problem of minimizing the expected risk $\mathbb{E}_{\xi\sim\mathsf{P}}[f(x;\xi)]$ with respect to $x\in\mathbb{R}^{d}$, where the random variable $\xi\in\mathcal{Z}\subseteq\mathbb{R}^{p}$ is distributed according to the unknown true data distribution $\mathsf{P}$. We approximate $\mathsf{P}$ with the empirical distribution $\widehat{\mathsf{P}}=\frac{1}{n}\sum_{i=1}^{n}\updelta_{\xi_{i}}$, where $\{\xi_{i}\}_{i\in\llbracket n\rrbracket}$ is the sample of size $n\in\mathbb{N}^{*}$, which leads to the empirical risk minimization problem:
$\operatorname*{minimize}_{x\in\mathbb{R}^{d}}\ F(x)\coloneqq\mathbb{E}_{\xi\sim\widehat{\mathsf{P}}}[f(x;\xi)]=\frac{1}{n}\sum_{i=1}^{n}f(x;\xi_{i}).$
When the sample size $n$ is large, the gradient of $F$ is approximated by its batch counterparts. Given a batch of samples $\mathcal{B}\subset\llbracket n\rrbracket$ of size $b\coloneqq|\mathcal{B}|$, we define the batch loss associated with the batch $\mathcal{B}$ by $F_{\mathcal{B}}(x)\coloneqq\frac{1}{b}\sum_{i\in\mathcal{B}}f(x;\xi_{i})$. If $f(\cdot;\xi)$ is continuously differentiable for any $\xi\in\mathcal{Z}$, the loss and batch loss gradients are given, respectively, by $\nabla F(x)=\frac{1}{n}\sum_{i=1}^{n}\nabla f(x;\xi_{i})$ and $\nabla F_{\mathcal{B}}(x)=\frac{1}{b}\sum_{i\in\mathcal{B}}\nabla f(x;\xi_{i})$. The batch gradient is an unbiased estimator for the loss function gradient, i.e., $\mathbb{E}_{\mathcal{B}}[\nabla F_{\mathcal{B}}(x)]=\nabla F(x)$ for any $x\in\mathbb{R}^{d}$. In many applications, including deep learning, the objective function $F$ is nonconvex. Thus, we consider the problem of finding a (first-order) $\varepsilon$-stationary point $x^{\star}\in\mathbb{R}^{d}$ that satisfies $\|\nabla F(x^{\star})\|^{2}\leqslant\varepsilon$, rather than a global minimum that is intractable.
4 Adaptive Sampling Methods
The adaptive batch schemes proposed in this paper are based on a family of adaptive sampling methods for stochastic optimization problems. Although such methods are extended to problems with different structures (see Section 2 for details), here we focus on the simplest unconstrained setting.
4.1 Norm Test
Byrd et al. (2012) proposed an adaptive sampling method, called the norm test or the norm condition, that was derived according to the following fundamental fact in stochastic optimization. If $f(\cdot;\xi)$ is a convex and continuously differentiable function for any $\xi\in\mathcal{Z}$, then $-\nabla F_{\mathcal{B}}(x)$ is a descent direction for $F$ at $x\in\mathbb{R}^{d}$ if there exists $\eta\in\left[0,1\right)$ such that
$\delta_{\mathcal{B}}(x)\coloneqq\|\nabla F_{\mathcal{B}}(x)-\nabla F(x)\|\leqslant\eta\|\nabla F(x)\|.$ (1)
See Appendix C for its proof. When the sample size $n$ is large, $\nabla F$ is not available, but $\delta_{\mathcal{B}}(x)$ can be approximated by
$\widehat{\delta}_{\mathcal{B}}(x)^{2}\coloneqq\frac{\mathrm{Var}_{i\in\mathcal{B}}(\nabla f(x;\xi_{i}))}{b}\cdot\frac{n-b}{n-1},$
where, for any vector-valued function $h\colon\mathbb{R}^{d}\times\mathcal{Z}\to\mathbb{R}^{d}$, the sample variance is defined by
$\mathrm{Var}_{i\in\mathcal{B}}(h(x;\xi_{i}))\coloneqq\frac{1}{b-1}\sum_{i\in\mathcal{B}}\left\lVert h(x;\xi_{i})-\mathbb{E}_{i\in\mathcal{B}}[h(x;\xi_{i})]\right\rVert^{2}.$ (2)
Therefore, as the sample size $n\to\infty$, condition (1) can be approximated as $\frac{1}{b}\mathrm{Var}_{i\in\mathcal{B}}(\nabla f(x;\xi_{i}))\leqslant\eta^{2}\|\nabla F_{\mathcal{B}}(x)\|^{2}$, which also corresponds to the case of sampling with replacement. In a practical implementation, the dynamic sample gradient algorithm with the norm test (Byrd et al., 2012) performs the following SGD update: $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$ with step size $\alpha_{k}>0$, and then checks the following (approximate) norm test condition:
$\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))}{b_{k}}\leqslant\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}.$ (3)
This can also be viewed as an approximation of the exact variance norm test in the stochastic setting:
$\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})-\nabla F(x_{k})\|^{2}\right]\leqslant\eta^{2}\|\nabla F(x_{k})\|^{2},$ (4)
meaning that the condition (1) holds in expectation. Here we abbreviate the conditional expectation on $\mathcal{F}_{k}$ (that is, the sigma algebra $\mathcal{F}_{k}\coloneqq\sigma(\{x_{1},\mathcal{B}_{1},\ldots,\mathcal{B}_{k-1}\})$) by $\mathbb{E}_{k}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$. Should the condition (3) not be met, the subsequent batch size is determined by the formula: $b_{k+1}=\left\lceil\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))/\left(\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right)\right\rceil$, and the batch $\mathcal{B}_{k+1}$ expands accordingly. The selection of the constant $\eta$ influences the probability of obtaining a descent direction, affecting the probability of increasing the batch size.
To ensure that the enlarged batch $\mathcal{B}_{k+1}$ satisfies the norm test condition, we employ a heuristic approach based on the assumption that the increase in batch size occurs gradually enough. Therefore, for any iterate $x_{k}$ produced by SGD, the following approximations hold: $\mathrm{Var}_{i\in\mathcal{B}_{k+1}}(\nabla f(x_{k};\xi_{i}))\approx\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))$ and $\|\nabla F_{\mathcal{B}_{k+1}}(x_{k})\|\approx\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|$, where $\mathcal{B}_{k}$ is the original batch and $\mathcal{B}_{k+1}$ is the new larger batch. In practice, these approximate equalities are not explicitly verified to maintain computational efficiency, and the batch for the next iteration, $\mathcal{B}_{k+1}$, is used directly without additional checks.
4.2 Inner Product Test
Bollapragada et al. (2018a) noted that the norm test often leads to a rapid increase in batch size, which could negate the benefits of adaptive sampling, and proposed the inner product test as an alternative. This test further controls the variance of the term $\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle$ to ensure that $-\nabla F_{\mathcal{B}_{k}}(x_{k})$ is a descent direction with high probability. The test selects a large enough batch size $b_{k}$ to meet the (exact variance) inner product test condition, which is satisfied when there is a $\vartheta>0$ ensuring that:
$\frac{1}{b_{k}}\mathbb{E}_{k}\left[\left(\langle\nabla f(x_{k};\xi_{i}),\nabla F(x_{k})\rangle-\|\nabla F(x_{k})\|^{2}\right)^{2}\right]\leqslant\vartheta^{2}\|\nabla F(x_{k})\|^{4}.$ (5)
As before, since $\nabla F$ is not available, the following (approximate) inner product test condition is usually used in practice:
$\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle)}{b_{k}}\leqslant\vartheta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{4},$ (6)
where the variance on the left hand side can be evaluated by (2). To establish a convergence rate, it is crucial to eliminate the possibility that $\nabla F_{\mathcal{B}_{k}}(x_{k})$ and $\nabla F(x_{k})$ are (near) orthogonal. Motivated by this need for theoretical guarantees, Bollapragada et al. (2018a) also introduce the following condition of the (exact variance) orthogonality test: there exists $\nu>0$ such that
$\frac{1}{b_{k}}\mathbb{E}_{k}\left[\left\lVert\nabla f(x_{k};\xi_{i})-\frac{\langle\nabla f(x_{k};\xi_{i}),\nabla F(x_{k})\rangle}{\|\nabla F(x_{k})\|^{2}}\nabla F(x_{k})\right\rVert^{2}\right]\leqslant\nu^{2}\|\nabla F(x_{k})\|^{2}.$ (7)
The above exact test also takes an approximate form for implementation purposes, given by
$\frac{1}{b_{k}}\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k};\xi_{i})-\frac{\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle}{\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\nabla F_{\mathcal{B}_{k}}(x_{k})\right)\leqslant\nu^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2},$ (8)
where the variance on the left hand side can again be computed via (2). These two tests are collectively referred to as the augmented inner product test.
Note that both (6) and (8) have to be satisfied, which collectively leads to the (approximate) augmented inner product test. If any of the conditions are violated, the batch size will be updated using the formula
$b_{k+1}=\left\lceil\max\left\{\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle)}{\vartheta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{4}},\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k};\xi_{i})-\frac{\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle}{\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\nabla F_{\mathcal{B}_{k}}(x_{k})\right)}{\nu^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\right\}\right\rceil,$
and the batch $\mathcal{B}_{k+1}$ is augmented with the above batch size. The constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ have to be chosen in practice. In implementation, we might also use the inner product test alone, since the (near) orthogonality of $\nabla F_{\mathcal{B}_{k}}(x_{k})$ and $\nabla F(x_{k})$ has not been observed in practice as mentioned in Bollapragada et al. (2018a), and the computation of the quantities in (8) would induce extra computational overhead for large-scale training. Then the batch size is updated using the formula $b_{k+1}=\left\lceil\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle)/\left(\vartheta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{4}\right)\right\rceil$.
4.3 Adaptive Sampling Methods for Adaptive Gradient Methods
We focus on two simple adaptive gradient methods, AdaGrad (Duchi et al., 2011; McMahan and Streeter, 2010) and AdaGrad-Norm (Streeter and McMahan, 2010), whose step sizes are computed adaptively based on the magnitudes of previous (batch) stochastic gradients. AdaGrad was proposed for online convex optimization, which takes the following update:
$(\forall k\in\mathbb{N}^{*})\quad v_{k}=v_{k-1}+g_{k}^{2},\quad x_{k+1}=x_{k}-\alpha g_{k}\odot v_{k}^{-\nicefrac{{1}}{{2}}},$ (9)
where $g_{k}\coloneqq\nabla F_{\mathcal{B}_{k}}(x_{k})$, $\alpha>0$ is a constant step size (or learning rate), $\odot$ denotes the Hadamard product, and the power operations are performed coordinate-wise. Since $(v_{k})_{k\in\mathbb{N}}\subset\mathbb{R}_{++}^{d}$ is a sequence of positive vectors, AdaGrad has adaptive coordinate-wise step sizes.
AdaGrad-Norm, also known as SGD with AdaGrad stepsizes, is the norm version of AdaGrad. It is given by
$(\forall k\in\mathbb{N}^{*})\quad v_{k}=v_{k-1}+\left\lVert g_{k}\right\rVert^{2},\quad x_{k+1}=x_{k}-\alpha g_{k}/\sqrt{v_{k}},$ (10)
where $(v_{k})_{k\in\mathbb{N}}\subset\mathbb{R}_{++}$ is a sequence of positive scalars. The scalar-valued adaptive step size $\alpha/\sqrt{v_{k}}$ makes the convergence of AdaGrad-Norm easier to analyze (Ward et al., 2019, 2020; Faw et al., 2022; Attia and Koren, 2023).
Although the norm test and the (augmented) inner product test were originally developed for SGD, their potential applicability to adaptive gradient methods is intriguing, given the widespread use of such methods in deep learning. Moreover, these tests focus on the approximation of loss gradients through batch gradients, making them optimizer-agnostic. As a result, adapting these batch size schemes for AdaGrad and AdaGrad-Norm—resulting in the creation of AdAdaGrad and AdAdaGrad-Norm, respectively—simply involves substituting the SGD updates with those from AdaGrad and AdaGrad-Norm. The complete pseudocode for these algorithms is provided in Algorithm 1.
The coordinate-wise nature of adaptive step sizes in AdaGrad, along with the simplicity of theoretical analysis, necessitates a coordinate-wise variant of AdAdaGrad. This variant introduces a more stringent coordinate-wise condition for the exact variance norm test: for each $j\in\llbracket d\rrbracket$,
$\mathbb{E}_{k}\left[\left(\partial_{j}F_{\mathcal{B}_{k}}(x_{k})-\partial_{j}F(x_{k})\right)^{2}\right]\leqslant\eta^{2}(\partial_{j}F(x_{k}))^{2},$ (11)
with its approximate form being $\frac{1}{b_{k}(b_{k}-1)}\sum_{i\in\mathcal{B}_{k}}\left(\partial_{j}f(x_{k};\xi_{i})-\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\leqslant\eta^{2}(\partial_{j}F_{\mathcal{B}_{k}}(x_{k}))^{2}$. However, the existence of inner products in the inner product test and the orthogonality test precludes the derivation of their coordinate-specific versions.
5 Convergence Analysis
We provide two sets of convergence rate results for AdAdaGrad and AdAdaGrad-Norm. We first establish the convergence rate results of the norm test and the augmented inner product test for AdaGrad-Norm and AdaGrad for smooth nonconvex objectives. Then, we provide a convergence rate analysis of AdAdaGrad-Norm for non-uniformly smooth objectives. Proofs of this section can be found in Appendix C. We might invoke the following assumptions in our analysis.
Assumption 1 ($L$-Lipschitz smoothness).
The function $F\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is continuously differentiable, bounded below by $F^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}F(x)=F(x^{\star})\in\mathbb{R}$ with $x^{\star}=\operatorname*{argmin}_{x\in\mathbb{R}^{d}}F(x)$, and $L$-Lipschitz smooth, i.e., its gradient $\nabla F$ is $L$-Lipschitz continuous with constant $L\geqslant 0$: for any $x,y\in\mathbb{R}^{d}$, $\|\nabla F(x)-\nabla F(y)\|\leqslant L\left\lVert x-y\right\rVert$.
[…]
Assumption 2 ($(L_{0},L_{1})$-smoothness).
The function $F$ is $(L_{0},L_{1})$-smooth with constants $(L_{0},L_{1})\in\mathbb{R}_{+}^{2}$, i.e., for any $x,y\in\mathbb{R}^{d}$, $\left\lVert\nabla F(x)-\nabla F(y)\right\rVert\leqslant(L_{0}+L_{1}\left\lVert\nabla F(x)\right\rVert)\left\lVert x-y\right\rVert$.
Aside from the smoothness assumption, the conditions of the two adaptive sampling methods imply the following assumption on the second moment of the stochastic gradient of (non)convex functions. Using the nomenclature in Khaled and Richtárik (2023), we introduced the expected strong growth condition (Vaswani et al., 2019) below.
Definition 1 (Expected strong growth).
For any batch $\mathcal{B}\subset\llbracket n\rrbracket$, the expected strong growth (E-SG) condition is given by $(\forall x\in\mathbb{R}^{d})\;\mathbb{E}[\|\nabla F_{\mathcal{B}}(x)\|^{2}]\leqslant\tau\|\nabla F(x)\|^{2}$, for some constant $\tau>0$. This condition can be relaxed in a sense that it only holds at each iteration, i.e.,
$(\forall k\in\mathbb{N}^{*})\quad\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant\tau\|\nabla F(x_{k})\|^{2},$ (12)
for some constant $\tau>0$ and $(x_{k})_{k\in\mathbb{N}^{*}}$ are the iterates generated by some stochastic gradient methods.
Proposition 5.1 (Informal).
For every iteration $k\in\mathbb{N}^{*}$, if the conditions of the exact variance norm test hold with constant $\eta\in\left(0,1\right)$ and the conditions of the exact variance augmented inner product test hold with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ respectively, then the E-SG condition (12) holds with constant $\tau>0$, where $\tau=1+\eta^{2}$ for the exact variance norm test and $\tau=1+\vartheta^{2}+\nu^{2}$ for the exact variance inner product test respectively.
A more precise statement and its proof are given in Appendix C. Although the expected strong growth condition holds in over-parameterized models (Vaswani et al., 2019) and is widely adopted in deep learning, the constant $\tau$ is usually problem-specific and unknown. In contrast, the two adaptive sampling methods impose a much weaker assumption; they require that this condition holds only for the iterates $(x_{k})_{k\in\mathbb{N}^{*}}$. In a practical implementation, the chosen constant $\tau$ is pre-specified and fixed, which is also likely to be smaller than the value for the global expected strong growth condition. If the (approximate) condition does not hold, the batch size is increased so that it holds for the next iterate (see Section 4).
5.1 Convergence Results
We first establish high probability convergence rates of the adaptive batch size schemes for AdaGrad-Norm and AdaGrad, substantially extending existing convergence rate results (in expectation) for SGD; see e.g., De et al. (2016, 2017); Bollapragada et al. (2018a). AdaGrad-Norm with the norm test or the augmented inner product test with any constant initial step size $\alpha>0$ enjoys the following high probability convergence bound for nonconvex functions.
Theorem 5.1 (AdAdaGrad-Norm).
Suppose that Assumption 1 holds. Let $(x_{k})_{k\in\llbracket K\rrbracket}$ be the AdaGrad-Norm iterates (10) with any step size $\alpha>0$, where the batch sizes $(b_{k})_{k\in\llbracket K\rrbracket}$ are chosen such that either the (exact variance) norm test (4) with constant $\eta\in(0,1)$ or the (exact variance) augmented inner product test (5) and (7) with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ is satisfied at each iteration. Then, with probability at least $1-\delta$, we have $\min_{k\in\llbracket K\rrbracket}\left\lVert\nabla F(x_{k})\right\rVert^{2}=\mathscr{O}\left(1/(K\delta^{2})\right)$.
To prove the convergence of AdaGrad which has coordinate-wise adaptive step sizes, we need the E-SG condition to hold coordinate-wise for the iterates as well. This indeed holds when we invoke a coordinate-wise version of the norm test, depicted in the following proposition.
Proposition 5.2 (Coordinate-wise expected strong growth).
For every iteration $k\in\mathbb{N}^{*}$, if the condition of the coordinate-wise exact variance norm test (11) holds with constant $\eta\in(0,1)$, then the coordinate-wise E-SG condition is satisfied at each iteration $k\in\mathbb{N}^{*}$, i.e., for all $(k,j)\in\mathbb{N}^{*}\times\llbracket d\rrbracket$, we have
$\mathbb{E}_{k}\left[\left(\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\right]\leqslant(1+\eta^{2})(\partial_{j}F(x_{k}))^{2}.$ (13)
Then, AdaGrad with the coordinate-wise norm test enjoys a similar high probability convergence guarantee.
Theorem 5.2 (AdAdaGrad).
Suppose that Assumption 1 holds. Let $(x_{k})_{k\in\llbracket K\rrbracket}$ be the AdaGrad iterates (9) with any step size $\alpha>0$, where the batch sizes $(b_{k})_{k\in\llbracket K\rrbracket}$ are chosen so that the coordinate-wise (exact variance) norm test (11) is satisfied at each iteration. Then, with probability at least $1-\delta$, we have $\min_{k\in\llbracket K\rrbracket}\left\lVert\nabla F(x_{k})\right\rVert^{2}=\mathscr{O}\left(1/(K\delta^{2})\right)$.
The above theorem establishes a sublinear convergence rate (with high probability) for AdAdaGrad for nonconvex functions, while such a rate (in expectation) has been established for SGD in Bollapragada et al. (2018a), Theorem 3.4.
Finally, relaxing the uniform (Lipschitz) smoothness assumption, we show that AdAdaGrad-Norm also converges for $(L_{0},L_{1})$-smooth functions, but required the initial step size to be upper bounded and hence the knowledge of $L_{1}$.
Theorem 5.3 ($(L_{0},L_{1})$-smooth AdAdaGrad-Norm).
Suppose that Assumption 2 holds. Let $(x_{k})_{k\in\llbracket K\rrbracket}$ be the AdaGrad-Norm iterates (10). Also suppose that the batch sizes $(b_{k})_{k\in\llbracket K\rrbracket}$ are chosen so that either the (exact variance) norm test (4) with constant $\eta\in(0,1)$ or the (exact variance) augmented inner product test (5) and (7) with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ is satisfied at each iteration. If the AdaGrad-Norm step size $\alpha\leqslant\min\{\omega/(2\rho_{1}\tau),\sqrt{\omega/(2\rho_{1}\tau)}\}/L_{1}$, where $(\rho_{1},\rho_{2},\omega)\in\mathbb{R}_{++}^{3}$ are constants such that $11/\rho_{1}+\rho_{1}/\rho_{2}+2\omega<$ 1 and $\tau\geqslant\min\{\omega/(2\min\{\rho_{1},\rho_{2}\}),1\}$, with $\tau=1+\eta^{2}$ for the norm test and $\tau=1+\vartheta^{2}+\nu^{2}$ for the augmented inner product test, then with probability at least $1-\delta$, we have $\min_{k\in\llbracket K\rrbracket}\left\lVert\nabla F(x_{k})\right\rVert^{2}=\mathscr{O}\left(1/(K\delta^{2})\right)$.
A similar convergence result can be derived for AdaGrad using a similar coordinate-wise argument for Theorem 5.2.
6 Numerical Experiments
We evaluate the performance of the norm test and the (augmented) inner product test with AdaGrad(-Norm) and SGD for image classification, employing logistic regression (Section D.1) and a three-layer CNN on the MNIST dataset (LeCun et al., 1998), and a three-layer CNN (Section D.3) and ResNet-18 (He et al., 2016) on the CIFAR-10 dataset (Krizhevsky, 2009). We note that training larger models often requires multiple workers and data parallelism, such as Distributed Data Parallel (DDP) (Li et al., 2020) and Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023). Extending and implementing our proposed schemes under data parallelism presents additional complexities and remains an area for future research. We thus concentrate on smaller models and datasets with the goal of demonstrating the concept rather than achieving state-of-the-art results. Given the typically large number of parameters $d$, conducting coordinate-wise norm tests for AdAdaGrad is not computationally practical, so the standard norm test is applied. Details of the experiments, such as the hyperparameter setting, and additional results can be found in Appendix D.
6.1 Three-layer CNN on MNIST
Table 1: Three-layer CNN on MNIST
Scheme test steps bsz. loss acc.
SGD N/A 366 16384 0.51 0.87
SGD N/A 183 32768 1.54 0.75
AdaGrad N/A 199 30000 0.10 0.97
AdaGrad N/A 149 40000 0.13 0.96
AdaSGD $\eta=0.10$ 256 23546 0.79 0.83
AdaSGD $\eta=0.25$ 383 15627 0.48 0.88
AdAdaGrad-Norm $\eta=0.10$ 226 26567 0.88 0.83
AdAdaGrad-Norm $\eta=0.25$ 435 13830 0.54 0.87
AdAdaGrad $\eta=0.10$ 149 40057 0.15 0.96
AdAdaGrad $\eta=0.25$ 198 30152 0.13 0.97
We first train a three-layer CNN on the MNIST classification problem with a nonconvex objective. Our experiments were carried out with an equal training budget of 6 million samples (equivalent to 100 epochs), setting a maximum batch size of 60,000 (i.e., the full batch) for all approaches. We can measure training efficiency based on the number of (gradient) steps, instead of wall-clock time (available in Table 5), which is device dependent (Shallue et al., 2019). Adaptive batch size methods begin from a small batch size 8 and gradually increase to the maximum possible batch size of 60000 (full batch). In Figure 1, AdAdaGrad outperforms AdAdaGrad-Norm and AdaSGD in terms of validation accuracy (generalization) by a certain margin. AdAdaGrad-Norm enjoys similar performance to AdaSGD, in spite of slightly higher training loss. We also observe from Table 1 that AdAdaGrad using the norm test with $\eta=0.1$ is able to achieve a validation accuracy of 96% with only 149 iterations with an average batch size of more than 40,000. It uses full batches for the last 70% of its training budget, taking full advantage of the available GPU memory. Referring to Table 5 in Section D.2, we observe the generalization gap between small constant batch sizes and large batch sizes, while using small batch sizes requires substantially longer training time and a higher number of steps, leading to lower training efficiency. Our proposed methods are capable of balancing this inevitable trade-off between training efficiency and generalization by introducing adaptive batch size schemes, without the need of extensive tuning of learning rates or pre-specified learning rate schedules.
Figure 1: Training loss, validation accuracy and batch sizes of AdaSGD, AdAdaGrad and AdAdaGrad-Norm for a three-layer CNN on the MNIST dataset.
6.2 ResNet-18 on CIFAR-10
We then train a ResNet-18 for image classification on the harder CIFAR-10 dataset. We use a training budget of 10 million samples (200 epochs) and a maximum batch size of 50,000 samples. Although the focus of this work is on AdaGrad, we also empirically study the effect of adaptive batch size schemes for Adam attributed to its ubiquity in deep learning, which is known as AdAdam.
AdAdaGrad.
In Figure 2, we observe that we need to choose a rather small $\eta$ to make use of full batches for training at later stages, when this larger model is trained. Comparing AdaGrad with a constant batch size of 50000 and AdAdaGrad with $\eta=0.025$ (see also Table 2), AdAdaGrad is able to use full batches in most of the later training stages, while achieving high accuracy with 23 more steps. In general (see also Table 10 in Section D.4),our proposed scheme is able to close the generalization gap between smaller and larger constant batches (e.g., the curves of $\eta=0.075$ and $0.1$ lying in the gap between those of constant batch sizes 4096 and 8192).
Figure 2: AdaGrad and AdAdaGrad for ResNet-18 on the CIFAR-10 dataset.
Figure 3: Adam and AdAdam for ResNet-18 on the CIFAR-10 dataset.
AdAdam.
In Figure 3, we observe a similar trend of AdAdam to AdAdaGrad but with faster convergence and larger batch size increases. It is worth noticing from Table 2 that AdAdam with $\eta=0.1$ and an average batch size of 8880 outperforms Adam with a smaller constant batch size of 8192 in terms of validation accuracy, while requiring almost 100 fewer steps. This indicates that our proposed scheme is even more beneficial when used with Adam.
Table 2: ResNet-18 on CIFAR-10
Scheme test steps bsz. loss acc.
AdaGrad N/A 305 32768 0.9684 0.5816
AdaGrad N/A 199 50000 1.3625 0.4708
Adam N/A 1220 8192 0.0004 0.8946
Adam N/A 610 16384 0.0028 0.8628
AdAdaGrad $\eta=0.025$ 222 44934 1.2770 0.5107
AdAdaGrad $\eta=0.05$ 485 20615 0.6204 0.7079
AdAdam $\eta=0.05$ 426 23463 0.0061 0.8228
AdAdam $\eta=0.1$ 1126 8880 0.0000 0.9042
7 Concluding Remarks
In this work, we demonstrate the versatility of adaptive sampling methods as generic adaptive batch size schemes suitable for adaptive gradient optimizers, supported by convergence guarantees and numerical results. This revelation opens up a myriad of intriguing research directions for adaptive batch size schemes as a promising and flexible option for large-scale model training. On the theoretical front, delving deeper into the convergence guarantees of this class of methods when combined with other stochastic gradient optimizers, such as momentum-based methods, as well as proximal SGD methods for constrained problems with deterministic nonsmooth regularizers, would be fascinating. Practically, exploring the implementation of adaptive batch size schemes under various parallelism paradigms for large-scale distributed training—including data, tensor, and pipeline parallelism (Shoeybi et al., 2019; Rajbhandari et al., 2020; Smith et al., 2022; Zhao et al., 2023)—is worthwhile. This exploration aims to ensure these schemes are viable for large-scale applications, such as the (pre-)training of autoregressive language and image models. Furthermore, examining the impact of adaptive batch size schemes for adaptive gradient methods, in contrast to those designed for SGD, particularly for language models based on transformers in addition to CNNs for vision tasks discussed in this paper, is crucial. This is because, unlike the marginal utility of adaptive methods for CNNs and RNNs (Wilson et al., 2017), adaptive gradient methods such as Adam significantly outperform SGD in optimizing transformers (Zhang et al., 2020b, c; Jiang et al., 2022; Kunstner et al., 2023; Pan and Li, 2023; Ahn et al., 2023).
Acknowledgments
The research of Han Liu is supported by NIH R01LM01372201, NSF RI 1840857, NSF TRIPOD 1740735, NSF DMS1454377-CAREER, NSF IIS 1546482, along with an Alfred P. Sloan Fellowship. The research of Mladen Kolar is supported in part by NSF ECCS-2216912. This research is supported in part through the computational resources and staff contributions provided for the Data Science Institute at the University of Chicago, through its AI + Science Research Funding Initiatives.
Appendix
Appendix A Further Related Work
A.1 Adaptive Sampling Methods
Adaptive sampling methods have also been extended to tackle problems outside the unconstrained optimization setting. Xie et al. (2023) proposed a proximal extension of the norm test and the inner product test to minimize a convex composite objective, which comprises a stochastic function and a deterministic, potentially nonsmooth function. Beiser et al. (2023) explored deterministic constrained problems, including those with nonconvex objectives. Moreover, adaptive sampling methods have been applied to various optimization problems and algorithms, such as (Bollapragada et al., 2018b) for L-BFGS, (Berahas et al., 2022) for sequential quadratic programming (SQP) in the context of equality-constrained problems, (Bollapragada et al., 2023) for augmented Lagrangian methods, and (Bollapragada and Wild, 2023) for quasi-Newton methods.
A.2 Convergence of Stochastic Gradient Methods
We give a brief overview on the convergence results of SGD and adaptive gradient methods. The convergence of SGD for smooth nonconvex functions was first analyzed in (Ghadimi and Lan, 2013), assuming a uniform bound on the variance of stochastic gradients. Arjevani et al. (2023) gave a tight lower bound. Bottou et al. (2018) extended the convergence result with the so-called affine variance noise model. Different assumptions on the second moment of stochastic gradients appearing in the literature were reviewed in (Khaled and Richtárik, 2023), with a view to developing a general convergence theory of SGD for smooth nonconvex functions.
Another line of work concerns the convergence guarantees in expectation of adaptive gradient methods. Ward et al. (2019, 2020); Li and Orabona (2019); Faw et al. (2022) each gave a rate of convergence of $\widetilde{\mathscr{O}}(1/\sqrt{K})$ for AdaGrad under different assumptions on stochastic gradients. Défossez et al. (2022) gave a simple proof on the convergence of both AdaGrad and a simplified version of Adam through a unified formulation.
Regarding high probability convergence bounds, Ghadimi and Lan (2013) established convergence with high probability for SGD with properly tuned learning rate, given knowledge of the smooth and sub-Gaussian stochastic gradient noise bounds. Under the same assumptions, Zhou et al. (2018); Li and Orabona (2020) also established similar results for (delayed) AdaGrad. Again, under different assumptions on stochastic gradients and proof techniques, Kavis et al. (2022); Faw et al. (2023); Wang et al. (2023); Attia and Koren (2023); Liu et al. (2023) obtained high probability convergence rates for AdaGrad(-Norm).
Appendix B The Overall Algorithm
Algorithm 1 Adaptive Batch Size Schemes for (Adaptive) Stochastic Gradient Methods (AdaSGD, AdAdaGrad and AdAdaGrad-Norm)
[…]
Sample an i.i.d. data batch $\mathcal{B}_{k}\subseteq\mathcal{B}$ uniformly from $\mathcal{B}$ of size $b_{k}$
Compute the batch gradient $\nabla F_{\mathcal{B}_{k}}(x_{k})\coloneqq\frac{1}{b_{k}}\sum_{i\in\mathcal{B}_{k}}\nabla f(x_{k};\xi_{i})$Ifnorm test
Compute the variance of per-sample gradients in the batch $\mathcal{B}_{k}$, $\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))$IfCoordinate-wise
Compute the coordinate-wise norm test statistics
$\mathsf{T}_{j}(x_{k};\mathcal{B}_{k},\eta)\coloneqq\frac{1}{b_{k}-1}\sum_{i\in\mathcal{B}_{k}}\left(\partial_{j}f(x_{k};\xi_{i})-\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\big{/}\left(\eta^{2}(\partial_{j}F_{\mathcal{B}_{k}}(x_{k}))^{2}\right),\;j\in\llbracket d\rrbracket$
Compute the aggregate coordinate-wise norm test statistic $\mathsf{T}=\max_{j\in\llbracket d\rrbracket}\mathsf{T}_{j}$Else
Compute the norm test statistic
$\hskip 56.9055pt\mathsf{T}\equiv\mathsf{T}(x_{k};\mathcal{B}_{k},\eta)\coloneqq\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))\big{/}\left(\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right)$
[…]
Compute the variance of the inner product between the batch per-sample gradients and the batch gradient $\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle)$
Compute the inner product test statistic
$\hskip 42.67912pt\mathsf{T}_{\mathrm{ip}}(x_{k};\mathcal{B}_{k},\vartheta)\coloneqq\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle)\big{/}\left(\vartheta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{4}\right)$
Compute the variance of the discrepancy of orthogonality between the batch per-sample gradients and the batch gradient
$\mathsf{V}(x_{k};\mathcal{B}_{k})\coloneqq\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k};\xi_{i})-\frac{\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle}{\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\nabla F_{\mathcal{B}_{k}}(x_{k})\right)$
Compute the orthogonality test statistic
$\hskip 71.13188pt\mathsf{T}_{\mathrm{ortho}}(x_{k};\mathcal{B}_{k},\nu)\coloneqq\mathsf{V}(x_{k};\mathcal{B}_{k})/\left(\nu^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right)$
Compute the augmented inner product test statistic $\mathsf{T}=\max\{\mathsf{T}_{\mathrm{ip}},\mathsf{T}_{\mathrm{ortho}}\}$EndIf
$b_{k+1}=\max\{\lceil\mathsf{T}\rceil,b_{k}\}$
$x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$triangleright AdaSGD
or
$v_{k}=v_{k-1}+\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}$ and $x_{k+1}=x_{k}-\alpha\nabla F_{\mathcal{B}_{k}}(x_{k})/\sqrt{v_{k}}$triangleright AdAdaGrad-Norm
or
$v_{k}=v_{k-1}+\nabla F_{\mathcal{B}_{k}}(x_{k})^{2}$ and $x_{k+1}=x_{k}-\alpha\nabla F_{\mathcal{B}_{k}}(x_{k})\odot v_{k}^{-\nicefrac{{1}}{{2}}}$triangleright AdAdaGrad
$i\leftarrow i+b_{k}$
$k\leftarrow k+1$EndWhile
[…]
Appendix C Proofs of Main Text
We provide the omitted proofs of the main text in this section.
C.1 Proof of Equation 1
Proof.
Note that $-\nabla F_{\mathcal{B}}(x)$ is a descent direction for $F$ if and only if $\langle\nabla F_{\mathcal{B}}(x),\nabla F(x)\rangle\geqslant 0$. Then we have, for $\eta\in\left[0,1\right)$, for any $x\in\mathbb{R}^{d}$,
$\|\nabla F_{\mathcal{B}}(x)-\nabla F(x)\|^{2}=\|\nabla F_{\mathcal{B}}(x)\|^{2}-2\langle\nabla F_{\mathcal{B}}(x),\nabla F(x)\rangle+\|\nabla F(x)\|^{2}\leqslant\eta^{2}\|\nabla F(x)\|^{2}\\ \iff-2\langle\nabla F_{\mathcal{B}}(x),\nabla F(x)\rangle\leqslant-(1-\eta^{2})\|\nabla F(x)\|^{2}-\|\nabla F_{\mathcal{B}}(x)\|^{2}\leqslant 0,$
which always holds when $-\nabla F_{\mathcal{B}}(x)$ is a descent direction for $F$. ∎
C.2 Preparatory Definitions, Propositions and Lemmas
We give various additional technical definitions, propositions and lemmas before giving the proofs of the theorems.
C.2.1 Formal Statements of Proposition 5.1
The formal version of Proposition 5.1 can be formally stated as the following two propositions.
Proposition C.1 (Exact variance norm test).
Suppose that the condition of the exact variance norm test (4) is satisfied for every iteration $k\in\mathbb{N}^{*}$, i.e., for some constant $\eta>0$,
$\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})-\nabla F(x_{k})\|^{2}\right]\leqslant\eta^{2}\|\nabla F(x_{k})\|^{2},$
where $\mathcal{B}_{k}$ is an i.i.d. batch at iteration $k$ and $x_{k}\in\mathbb{R}^{d}$ is the iterate of iteration $k$ generated by a stochastic gradient method. Then, we have
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant(1+\eta^{2})\|\nabla F(x_{k})\|^{2}.$ (C.1)
Proof.
Notice that we have
$\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})-\nabla F(x_{k})\|^{2}\right]$ $=\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}-2\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle+\|\nabla F(x_{k})\|^{2}\right]$
$=\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right]-2\langle\mathbb{E}_{k}\left[\nabla F_{\mathcal{B}_{k}}(x_{k})\right],\nabla F(x_{k})\rangle+\|\nabla F(x_{k})\|^{2}$
$=\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right]-\|\nabla F(x_{k})\|^{2}.$
The condition of the exact variance norm test (4) implies
$\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right]-\|\nabla F(x_{k})\|^{2}\leqslant\eta^{2}\|\nabla F(x_{k})\|^{2}\implies\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant(1+\eta^{2})\|\nabla F(x_{k})\|^{2}.$
Proposition C.2 (Exact variance inner product test and orthogonality test).
Suppose that the conditions of the exact variance inner product test (5) and the exact variance orthogonality test (7) are satisfied for every iteration $k\in\mathbb{N}^{*}$, i.e., for some constant , i.e., for some constant $\vartheta>0$,
$\frac{1}{b_{k}}\mathbb{E}_{k}\left[\left(\langle\nabla\ell(x_{k};\xi_{i}),\nabla F(x_{k})\rangle-\|\nabla F(x_{k})\|^{2}\right)^{2}\right]\leqslant\vartheta^{2}\|\nabla F(x_{k})\|^{4},$
and for some constant $\nu>0$,
$\frac{1}{b_{k}}\mathbb{E}_{k}\left[\left\lVert\nabla\ell(x_{k};\xi_{i})-\frac{\langle\nabla\ell(x_{k};\xi_{i}),\nabla F(x_{k})\rangle}{\|\nabla F(x_{k})\|^{2}}\nabla F(x_{k})\right\rVert^{2}\right]\leqslant\nu^{2}\|\nabla F(x_{k})\|^{2},$
where $\mathcal{B}_{k}$ is an i.i.d. batch at iteration $k$ and $x_{k}\in\mathbb{R}^{d}$ is the iterate of iteration $k$ generated by a stochastic gradient method. Then, we have
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant(1+\vartheta^{2}+\nu^{2})\|\nabla F(x_{k})\|^{2}.$ (C.2)
Proof.
This is part of the result of (Bollapragada et al., 2018a), Lemma 3.1. We include its proof here with our notation for completeness.
Since (7) is satisfied, the following inequality holds:
$\mathbb{E}_{k}\left[\left\lVert\nabla F_{\mathcal{B}_{k}}(x_{k})-\frac{\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle}{\|\nabla F(x_{k})\|^{2}}\nabla F(x_{k})\right\rVert^{2}\right]\leqslant\nu^{2}\|\nabla F(x_{k})\|^{2}.$
Then, applying $\mathbb{E}_{k}[\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle]=\|\nabla F(x_{k})\|^{2}$, we have
$\;\quad\mathbb{E}_{k}\left[\left\lVert\nabla F_{\mathcal{B}_{k}}(x_{k})-\frac{\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle}{\|\nabla F(x_{k})\|^{2}}\nabla F(x_{k})\right\rVert^{2}\right]$
$=\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]-2\frac{\mathbb{E}_{k}\left[\left(\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle\right)^{2}\right]}{\|\nabla F(x_{k})\|^{2}}+\frac{\mathbb{E}_{k}\left[\left(\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle\right)^{2}\right]}{\|\nabla F(x_{k})\|^{2}}$
$=\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]-\frac{\mathbb{E}_{k}\left[\left(\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle\right)^{2}\right]}{\|\nabla F(x_{k})\|^{2}}$
$\leqslant\nu^{2}\|\nabla F(x_{k})\|^{2},$
which leads to
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant\frac{\mathbb{E}_{k}\left[\left(\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle\right)^{2}\right]}{\|\nabla F(x_{k})\|^{2}}+\nu^{2}\|\nabla F(x_{k})\|^{2}.$ (C.3)
It remains to bound the first term on the right hand side of (C.3), using the condition of the exact variance inner product test. Since (5) is satisfied, the following inequality holds:
$\mathbb{E}_{k}\left[\left(\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle-\|\nabla F(x_{k})\|^{2}\right)^{2}\right]\leqslant\vartheta^{2}\|\nabla F(x_{k})\|^{4}.$
This implies
$\mathbb{E}_{k}\left[\left(\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle\right)^{2}\right]\leqslant\|\nabla F(x_{k})\|^{4}+\vartheta^{2}\|\nabla F(x_{k})\|^{4}=(1+\vartheta^{2})\|\nabla F(x_{k})\|^{4}.$
Along with (C.3), this yields
$\mathbb{E}_{k}[\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}]\leqslant(1+\vartheta^{2}+\nu^{2})\|\nabla F(x_{k})\|^{2}.$
C.2.2 Proof of Proposition 5.2
We also give the proof of Proposition 5.2.
Proof.
Summing (13) over $j$ from $1$ to $d$ yields
$\sum_{i=1}^{d}\mathbb{E}_{k}\left[\left(\partial_{j}F_{\mathcal{B}_{k}}(x_{k})-\partial_{j}F(x_{k})\right)^{2}\right]$ $\leqslant\eta^{2}\sum_{i=1}^{d}(\partial_{j}F(x_{k}))^{2}$
$\mathbb{E}_{k}\left[\sum_{i=1}^{d}\left(\partial_{j}F_{\mathcal{B}_{k}}(x_{k})-\partial_{j}F(x_{k})\right)^{2}\right]$ $\leqslant\eta^{2}\sum_{i=1}^{d}(\partial_{j}F(x_{k}))^{2}$
$\mathbb{E}_{k}\left[\|\nabla F_{\mathcal{B}_{k}}(x_{k})-\nabla F(x_{k})\|^{2}\right]$ $\leqslant\eta^{2}\left\lVert\nabla F(x_{k})\right\rVert^{2},$
which is identical to (4). The desired conclusion follows from Proposition C.1. ∎
C.2.3 Technical Lemmas
We give a technical lemma concerning upper bounds for finite sums of sequences of nonnegative real numbers.
Lemma C.1.
Let $(a_{k})_{k\in\mathbb{N}}\subset\mathbb{R}_{+}$ be a sequence of nonnegative real numbers with $a_{0}>0$. Then, the following inequalities hold:
$\sum_{k=1}^{K}\frac{a_{k}}{\left(\sum_{i=0}^{k}a_{i}\right)^{\negthickspace\nicefrac{{3}}{{2}}}}$ $\leqslant\frac{2}{\sqrt{a_{0}}},$
$\sum_{k=1}^{K}\frac{a_{k}}{\sum_{i=0}^{k}a_{i}}$ $\leqslant\log\sum_{k=0}^{K}a_{k}-\log a_{0}.$
Proof.
The first inequality is due to
$\sum_{k=1}^{K}\frac{a_{k}}{\left(\sum_{i=0}^{k}a_{i}\right)^{\negthickspace\nicefrac{{3}}{{2}}}}\leqslant\int_{a_{0}}^{\sum_{k=0}^{K}a_{k}}\frac{1}{t^{\nicefrac{{3}}{{2}}}}\,\mathrm{d}t,$
whereas the second inequality is due to
$\sum_{k=1}^{K}\frac{a_{k}}{\sum_{i=0}^{k}a_{i}}\leqslant\int_{a_{0}}^{\sum_{k=0}^{K}a_{k}}\frac{1}{t}\,\mathrm{d}t.$
We state without proof a similar descent lemma for $(L_{0},L_{1})$-smooth functions. Its proof can be found in (Zhang et al., 2020a), Lemma A.3.
Lemma C.2 (Descent lemma for $(L_{0},L_{1})$-smooth functions).
If Assumption 2 is satisfied, then for $x,y\in\mathbb{R}^{d}$ satisfying $\left\lVert x-y\right\rVert\leqslant 1/L$,
$F(y)\leqslant F(x)+\langle\nabla F(x),y-x\rangle+\frac{L_{0}+L_{1}\left\lVert\nabla F(x)\right\rVert}{2}\left\lVert x-y\right\rVert^{2}.$
C.3 Convergence Results for AdaGrad-Norm and AdaGrad
C.3.1 Full Statement of Theorem 5.1
Theorem C.1 (AdAdaGrad-Norm).
Suppose that Assumption 1 holds. Let $(x_{k})_{k\in\mathbb{N}^{*}}$ be the AdaGrad-Norm iterates (10) with any step size $\alpha>0$, where the batch sizes $(b_{k})_{k\in\mathbb{N}^{*}}$ are chosen such that either the (exact variance) norm test (4) with constant $\eta\in(0,1)$ or the (exact variance) augmented inner product test (5) and (7) with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ is satisfied at each iteration. Then, with probability at least $1-\delta$, we have
$\min_{k\in\llbracket K\rrbracket}\left\lVert\nabla F(x_{k})\right\rVert^{2}\leqslant\frac{c_{3}(c_{1}+2c_{2}\log c_{3})}{K\delta^{2}}.$
The constants $c_{1}$, $c_{2}$ and $c_{3}$ are defined through $c_{1}\coloneqq\frac{2}{\alpha(1-\rho^{-1})}(F(x_{1})-F^{\star}+\tau\alpha\left\lVert\nabla F(x_{0})\right\rVert^{2}/(2\sqrt{v_{0}})+\tau L^{2}\alpha^{3}(1+\rho\tau)/\sqrt{v_{0}}-L\alpha^{2}\log(v_{0})/2)$, $c_{2}\coloneqq 2L\alpha/(1-\rho^{-1})$, and $c_{3}\coloneqq 2c_{1}\tau+8c_{2}\tau\log(c_{2}\tau+1)$, where $\rho\in\left(1,\infty\right)$ is a constant, with $\tau=1+\eta^{2}$ for the norm test and $\tau=1+\vartheta^{2}+\nu^{2}$ for the augmented inner product test.
C.3.2 Proof of Theorem 5.1
Proof.
By Assumption 1, we have
$F(x_{k+1})$ $\leqslant F(x_{k})+\langle\nabla F(x_{k}),x_{k+1}-x_{k}\rangle+\frac{L}{2}\left\lVert x_{k+1}-x_{k}\right\rVert^{2}$
$=F(x_{k})-\alpha\left\langle\nabla F(x_{k}),\frac{g_{k}}{\sqrt{v_{k}}}\right\rangle+\frac{L\alpha^{2}}{2}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$
Taking conditional expectation with respect to $\mathcal{F}_{k}$, we have
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$
The above inner product (first-order) term can be decomposed into a sum of two terms, a first order main term and an error term:
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle$ $=\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k-1}}}\right]\right\rangle+\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k}}}-\frac{1}{\sqrt{v_{k-1}}}\right)g_{k}\right]\right\rangle$
$=\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k}}}-\frac{1}{\sqrt{v_{k-1}}}\right)g_{k}\right]\right\rangle,$ (C.4)
which leads to
$\mathbb{E}_{k}[F(x_{k+1})]\\ \leqslant F(x_{k})-\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$ (C.5)
To upper bound the error term, we first prove a preparatory lemma.
Lemma C.1.
Define the auxiliary function $\varphi_{k}\coloneqq\left\lVert\nabla F(x_{k})\right\rVert^{2}/\sqrt{v_{k}}$ for $k\in\mathbb{N}$. The error term has the upper bound:
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle\\ \leqslant\frac{1}{2}\left(1+\frac{1}{\rho}\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\tau}{2}\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]+\frac{\tau L^{2}\alpha^{2}}{2}(1+\rho\tau)\frac{\left\lVert g_{k}\right\rVert^{2}}{v_{k+1}^{\nicefrac{{3}}{{2}}}}.$ (C.6)
Proof of Lemma C.1.
Note that
$\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}=\frac{v_{k}-v_{k-1}}{\sqrt{v_{k-1}}\sqrt{v_{k}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})}=\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k-1}}\sqrt{v_{k}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})}.$ (C.7)
Then, we have
$\quad\;\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle$
$=\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k-1}}\sqrt{v_{k}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})}g_{k}\right]\right\rangle$
$\leqslant\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{3}}{\sqrt{v_{k}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})}\right]$ by Cauchy–Schwarz’s inequality
$\leqslant\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right]$ since $v_{k}\geqslant\left\lVert g_{k}\right\rVert^{2}$
$\leqslant\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{2\sqrt{v_{k-1}}}+\frac{1}{2\sqrt{v_{k-1}}}\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right]^{2}$ by the A.M.-G.M. inequality
$\leqslant\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{2\sqrt{v_{k-1}}}+\frac{1}{2\sqrt{v_{k-1}}}\mathbb{E}_{k}[\left\lVert g_{k}\right\rVert^{2}]\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{(\sqrt{v_{k}}+\sqrt{v_{k-1}})^{2}}\right]^{2}$ by Hölder’s inequality
[…]
Since by (C.7),
$\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k-1}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})^{2}}\leqslant\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k-1}}\sqrt{v_{k}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})}=\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}},$
we have
$\frac{\tau\left\lVert\nabla F(x_{k})\right\rVert^{2}}{2\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{(\sqrt{v_{k}}+\sqrt{v_{k-1}})^{2}}\right]^{2}\leqslant\frac{\tau}{2}\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right].$
Furthermore, we can decompose the term
$\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]=\mathbb{E}_{k}\left[\frac{\left\lVert\nabla F(x_{k-1})\right\rVert^{2}}{\sqrt{v_{k-1}}}-\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k}}}\right]+\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}-\left\lVert\nabla F(x_{k-1})\right\rVert^{2}}{\sqrt{v_{k-1}}}.$
We then derive an upper bound for the term $\left\lVert\nabla F(x_{k})\right\rVert^{2}-\left\lVert\nabla F(x_{k-1})\right\rVert^{2}$. By the reverse triangle inequality and Assumption 1, we have
$\left\lVert\nabla F(x_{k})\right\rVert-\left\lVert\nabla F(x_{k-1})\right\rVert\leqslant\left\lVert\nabla F(x_{k})-\nabla F(x_{k-1})\right\rVert\leqslant L\left\lVert x_{k}-x_{k-1}\right\rVert.$
On the other hand, by the triangle inequality and Assumption 1, we have
$\left\lVert\nabla F(x_{k-1})\right\rVert\leqslant\left\lVert\nabla F(x_{k-1})-\nabla F(x_{k})\right\rVert+\left\lVert\nabla F(x_{k})\right\rVert\leqslant L\left\lVert x_{k}-x_{k-1}\right\rVert+\left\lVert\nabla F(x_{k})\right\rVert.$
Therefore, we obtain
$\left\lVert\nabla F(x_{k})\right\rVert^{2}-\left\lVert\nabla F(x_{k-1})\right\rVert^{2}$ $=\left(\left\lVert\nabla F(x_{k})\right\rVert-\left\lVert\nabla F(x_{k-1})\right\rVert\right)\left(\left\lVert\nabla F(x_{k})\right\rVert+\left\lVert\nabla F(x_{k-1})\right\rVert\right)$
$\leqslant L\left\lVert x_{k}-x_{k-1}\right\rVert(L\left\lVert x_{k}-x_{k-1}\right\rVert+2\left\lVert\nabla F(x_{k})\right\rVert)$
$=L^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+2L\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert.$
Hence,
$\quad\;\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]$
$\leqslant\mathbb{E}_{k}\left[\frac{\left\lVert\nabla F(x_{k-1})\right\rVert^{2}}{\sqrt{v_{k-1}}}-\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k}}}\right]+\frac{1}{\sqrt{v_{k-1}}}\left(L^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+2L\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert\right)$
$=\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]+\frac{1}{\sqrt{v_{k-1}}}\left(\frac{L^{2}\alpha^{2}}{v_{k-1}}\left\lVert g_{k-1}\right\rVert^{2}+\frac{2L\alpha}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert\left\lVert g_{k-1}\right\rVert\right)$
By the A.M.-G.M. inequality with $\rho>0$, we have
$\frac{2\tau L\alpha}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert\left\lVert g_{k-1}\right\rVert\leqslant\frac{\rho\tau^{2}L^{2}\alpha^{2}}{v_{k-1}}\left\lVert g_{k-1}\right\rVert^{2}+\frac{1}{\rho}\left\lVert\nabla F(x_{k})\right\rVert^{2}.$
Therefore, we get
$\quad\;\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle$
$\leqslant\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{2\sqrt{v_{k-1}}}+\frac{\tau}{2}\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]+\frac{\rho\tau^{2}L^{2}\alpha^{2}}{2}\cdot\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}+\frac{1}{2\rho}\cdot\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\tau L^{2}\alpha^{2}}{2}\cdot\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}$
$=\frac{1}{2}\left(1+\frac{1}{\rho}\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\tau}{2}\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]+\frac{\tau L^{2}\alpha^{2}}{2}(1+\rho\tau)\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}.$
We now continue with the proof of Theorem 5.1. Using Lemma C.1, (C.5) is upper bounded by
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\alpha}{2}\left(1+\frac{1}{\rho}\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\tau\alpha}{2}\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]\\ +\frac{\tau L^{2}\alpha^{3}}{2}(1+\rho\tau)\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$
Taking expectation with respect to $\mathcal{F}_{k}$ gives
$\mathbb{E}[F(x_{k+1})]\leqslant\mathbb{E}[F(x_{k})]-\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}-\frac{\alpha}{2}\left(1-\frac{1}{\rho}\right)\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]+\frac{\tau\alpha}{2}\mathbb{E}[\varphi_{k-1}-\varphi_{k}]\\ +\frac{L\alpha^{2}}{2}\mathbb{E}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}+\frac{\tau L^{2}\alpha^{3}}{2}(1+\rho\tau)\mathbb{E}\left[\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}\right].$ (C.8)
Summing the last two terms on the right hand side over $k$ from $1$ to $K$ and using Lemma C.1 yields
$\frac{L\alpha^{2}}{2}\sum_{k=1}^{K}\mathbb{E}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}+\frac{\tau L^{2}\alpha^{3}}{2}(1+\rho\tau)\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}\right]\\ \leqslant\frac{L\alpha^{2}}{2}(\mathbb{E}\log v_{K}-\log v_{0})+\tau L^{2}\alpha^{3}(1+\rho\tau)\frac{1}{\sqrt{v_{0}}}.$
Thus, summing (C.8) over $k$ from $1$ to $K$ and rearranging yields
$\frac{\alpha}{2}\left(1-\frac{1}{\rho}\right)\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\leqslant F(x_{1})-\mathbb{E}[F(x_{K+1})]+\frac{\tau\alpha}{2}\mathbb{E}[\varphi_{0}-\varphi_{K}]\\ +\frac{L\alpha^{2}}{2}(\mathbb{E}\log v_{K}-\log v_{0})+\tau L^{2}\alpha^{3}(1+\rho\tau)\frac{1}{\sqrt{v_{0}}},$
which implies
$\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\leqslant c_{1}+c_{2}\mathbb{E}\log v_{K},$
where $c_{1}$ and $c_{2}$ are defined through
$c_{1}$ $\coloneqq\frac{2}{\alpha(1-\rho^{-1})}\left(F(x_{1})-F^{\star}+\frac{\tau\alpha}{2}\cdot\frac{\left\lVert\nabla F(x_{0})\right\rVert^{2}}{\sqrt{v_{0}}}++\tau L^{2}\alpha^{3}(1+\rho\tau)\frac{1}{\sqrt{v_{0}}}-\frac{L\alpha^{2}}{2}\log v_{0}\right),$
$c_{2}$ $\coloneqq\frac{2L\alpha}{1-\rho^{-1}}.$
By Proposition 5.1, we also have
$c_{1}+c_{2}\mathbb{E}\log v_{K}\geqslant\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\geqslant\frac{1}{\tau}\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\geqslant\frac{1}{\tau}\mathbb{E}\left[\frac{1}{\sqrt{v_{K}}}\sum_{k=1}^{K}\left\lVert g_{k}\right\rVert^{2}\right],$
implying that, together with Jensen’s inequality, we have
$\frac{1}{\tau}\mathbb{E}[\sqrt{v_{K}}]=\frac{1}{\tau}\mathbb{E}\left[\frac{1}{\sqrt{v_{K}}}\sum_{k=1}^{K}\left\lVert g_{k}\right\rVert^{2}\right]\leqslant c_{1}+c_{2}\mathbb{E}\log v_{K}\leqslant c_{1}+2c_{2}\log\mathbb{E}[\sqrt{v_{K}}].$
We proceed by solving the above inequality for $\mathbb{E}[\sqrt{v_{k}}]$, with the following lemma.
Lemma C.2.
For any positive constants $(a,b)\in\mathbb{R}_{++}^{2}$, if $x\leqslant a+b\log x$ for any $x>0$, then $x\leqslant 2a-2b+4b\log(b/2+1)\leqslant 2a+4b\log(b/2+1)$.
Proof of Lemma C.2.
Let $h(x)\coloneqq x/2-\log x$. Then $h^{\prime}(x)=1/2-b/x=0$ if and only if $x=2b$. Since $h^{\prime\prime}(x)=b/x^{2}>0$ for any $x>0$, so $h$ achieves its minimum at $2b$, which means $h(x)\geqslant h(2b)=b-b\log(2b)$. By the A.M.-G.M. inequality, we have $b+2\geqslant 2\sqrt{2b}$ and taking logarithm yields $\log(b+2)\geqslant\log 2+\frac{1}{2}(\log 2+\log b)$. Therefore, we obtain $b\log(2b)=b(\log 2+\log b)\leqslant 2b(\log(b+2)-\log 2)=2b\log\left(b/2+1\right)$. Consequently, we get $x-b\log x=x/2+x/2-b\log x\geqslant x/2+b-b\log(2b)\geqslant x/2+b-2b\log\left(b/2+1\right)$. Therefore, we have $x/2+b-2b\log\left(b/2+1\right)\leqslant a$, which implies the desired inequality. ∎
By Lemma C.2, we have $\mathbb{E}[\sqrt{v_{k}}]\leqslant 2c_{1}\tau+8c_{2}\tau\log(c_{2}\tau+1)\eqqcolon c_{3}$.
Since, by Hölder’s inequality,
$c_{1}+2c_{2}\log\mathbb{E}[\sqrt{v_{k}}]\geqslant\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\\ \geqslant\mathbb{E}\left[\frac{1}{\sqrt{v_{K}}}\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}\right]\geqslant\frac{\mathbb{E}\left[\sqrt{\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}}\right]^{2}}{\mathbb{E}[\sqrt{v_{K}}]},$
we have
$\mathbb{E}\left[\sqrt{\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}}\right]^{\negthickspace 2}\leqslant\mathbb{E}[\sqrt{v_{K}}]\cdot\left(c_{1}+2c_{2}\log\mathbb{E}[\sqrt{v_{K}}]\right)\leqslant c_{3}(c_{1}+2c_{2}\log c_{3}).$
Then by Markov’s inequality, we obtained the desired result. ∎
C.3.3 Proof of Theorem 5.2
Proof.
Following similar arguments in the proof of Theorem 5.1, using Assumption 1 and taking conditional expectation with respect to $\mathcal{F}_{k}$, we have
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\left\langle\nabla F(x_{k}),\frac{1}{\sqrt{v_{k-1}}}\odot\nabla F(x_{k})\right\rangle\\ +\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)\odot g_{k}\right]\right\rangle+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{1}{\sqrt{v_{k}}}\odot g_{k}\right\rVert^{2}.$ (C.9)
The inner product can be written as a sum:
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)\odot g_{k}\right]\right\rangle=\sum_{i=1}^{d}\mathbb{E}_{k}\left[\partial_{j}F(x_{k})\left(\frac{1}{\sqrt{v_{k-1,j}}}-\frac{1}{\sqrt{v_{k,j}}}\right)g_{k,j}\right],$
where $g_{k,j}\coloneqq[g_{k}]_{j}=\partial_{j}F_{\mathcal{B}_{k}}(x_{k})$. Note that
$\quad\;\mathbb{E}_{k}\left[\partial_{j}F(x_{k})\left(\frac{1}{\sqrt{v_{k-1,j}}}-\frac{1}{\sqrt{v_{k,j}}}\right)g_{k,j}\right]$
$\leqslant\frac{(\partial_{j}F(x_{k}))^{2}}{2\sqrt{v_{k-1,j}}}+\frac{1}{2\sqrt{v_{k,j}}}\left(\mathbb{E}_{k}\left[\frac{g_{k,j}^{2}}{\sqrt{v_{k,j}}+\sqrt{v_{k-1,j}}}\right]\right)^{\negthickspace 2}$
$\leqslant\frac{(\partial_{j}F(x_{k}))^{2}}{2\sqrt{v_{k-1,j}}}+\frac{1}{2\sqrt{v_{k,j}}}\mathbb{E}_{k}[g_{k,j}^{2}]\cdot\mathbb{E}_{k}\left[\frac{g_{k,j}^{2}}{(\sqrt{v_{k,j}}+\sqrt{v_{k-1,j}})^{2}}\right]$ by Hölder’s inequality
$\leqslant\frac{(\partial_{j}F(x_{k}))^{2}}{2\sqrt{v_{k-1,j}}}+\frac{\tau}{2\sqrt{v_{k,j}}}(\partial_{j}F(x_{k}))^{2}\cdot\mathbb{E}_{k}\left[\frac{g_{k,j}^{2}}{(\sqrt{v_{k,j}}+\sqrt{v_{k-1,j}})^{2}}\right]$ by Proposition 5.2.
Now we define a coordinate-wise auxiliary function $\widetilde{\varphi}_{k,j}\coloneqq(\partial_{j}F(x_{k}))^{2}/\sqrt{v_{k,j}}$. Then, similar to the proof of Theorem 5.1, we get
$\mathbb{E}_{k}\left[\partial_{j}F(x_{k})\left(\frac{1}{\sqrt{v_{k-1,j}}}-\frac{1}{\sqrt{v_{k,j}}}\right)g_{k,j}\right]\leqslant\frac{1}{2}\left(1+\frac{1}{\rho}\right)\frac{(\partial_{j}F(x_{k}))^{2}}{\sqrt{v_{k-1,j}}}+\frac{\tau}{2}\mathbb{E}_{k}[\widetilde{\varphi}_{k-1,j}-\widetilde{\varphi}_{k,j}]\\ +\frac{\tau L^{2}\alpha^{2}}{2}(1+\rho\tau)\frac{1}{\sqrt{v_{k-1,j}}}\left\lVert\frac{1}{\sqrt{v_{k-1}}}\odot g_{k-1}\right\rVert^{2}.$ (C.10)
Using (C.10) and summing (C.9) over $k$ from $1$ to $K$, we get
$\frac{\alpha}{2}\left(1-\frac{1}{\rho}\right)\sum_{k=1}^{K}\sum_{i=1}^{d}\mathbb{E}\frac{(\partial_{j}F(x_{k}))^{2}}{\sqrt{v_{k-1,j}}}\leqslant F(x_{1})-\mathbb{E}[F(x_{K+1})]+\frac{\tau\alpha}{2}\sum_{i=1}^{d}\mathbb{E}[\widetilde{\varphi}_{0,j}-\widetilde{\varphi}_{K,j}]\\ +\left(\frac{\tau L^{2}\alpha^{3}(1+\rho\tau)}{\min_{j\in\llbracket d\rrbracket}\sqrt{v_{0,j}}}+\frac{L\alpha^{2}}{2}\right)\sum_{i=1}^{d}\left(\mathbb{E}\log v_{K,j}-\log v_{0,j}\right).$
Now we define
$\widetilde{c}_{2}$ $\coloneqq\frac{2}{1-\rho^{-1}}\left(\frac{\tau L^{2}\alpha^{2}(1+\rho\tau)}{\min_{j\in\llbracket d\rrbracket}\sqrt{v_{0,j}}}+\frac{L\alpha}{2}\right),$
$\widetilde{c}_{1}$ $\coloneqq\frac{2}{\alpha(1-\rho^{-1})}(F(x_{1})-F^{\star})+\frac{\tau}{1-\rho^{-1}}\sum_{i=1}^{d}\mathbb{E}[\widetilde{\varphi}_{0,j}-\widetilde{\varphi}_{K,j}]-\widetilde{c}_{2}\sum_{i=1}^{d}\log v_{0,j}.$
Following the proof of Theorem 5.1, we also have
$\frac{1}{\tau}\mathbb{E}\left[\sum_{i=1}^{d}\sqrt{v_{K,j}}\right]\leqslant\widetilde{c}_{1}+\widetilde{c}_{2}\mathbb{E}\left[\sum_{i=1}^{d}\log v_{K,j}\right]\leqslant\widetilde{c}_{1}+2d\widetilde{c}_{2}\log\left(\mathbb{E}\left[\sum_{i=1}^{d}\sqrt{v_{K,j}}\right]\right),$
by Jensen’s inequality. Then, by Lemma C.2, we have the following inequality
$\mathbb{E}\left[\sum_{i=1}^{d}\sqrt{v_{K,j}}\right]\leqslant 2\widetilde{c}_{1}\tau+8d\widetilde{c}_{2}\tau\log(d\widetilde{c}_{2}\tau+1).$
Again, by Markov’s inequality, the desired high probability convergence result follows. ∎
C.3.4 Proof of Theorem 5.3
Proof.
As will be seen later, under the choice of the parameters $(\rho_{1},\rho_{2},\omega)$, the constant step size satisfies $\alpha\leqslant 1/L_{1}$ and $\left\lVert\frac{\nabla F(x_{k})}{\sqrt{v_{k}}}\right\rVert\leqslant 1$, we have
$\left\lVert x_{k+1}-x_{k}\right\rVert=\alpha\left\lVert\frac{\nabla F(x_{k})}{\sqrt{v_{k}}}\right\rVert\leqslant 1/L_{1}.$
By Lemma C.2 and taking conditional expectation with respect to $\mathcal{F}_{k}$, we have
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\alpha^{2}\frac{L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$ (C.11)
Let us recall from (C.4) that we have
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle=\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k}}}-\frac{1}{\sqrt{v_{k-1}}}\right)g_{k}\right]\right\rangle.$ (C.12)
The second order term in (C.11) is equal to
$\frac{L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}=\frac{L_{0}\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}+\frac{L_{1}\alpha^{2}\left\lVert\nabla F(x_{k})\right\rVert}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$
Also, following the proof of Theorem 5.1, the second term in (C.12) can be upper bounded by
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k}}}-\frac{1}{\sqrt{v_{k-1}}}\right)g_{k}\right]\right\rangle\leqslant\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right].$
Since $v_{k}$ is nondecreasing in $k$, we have
$\frac{L_{1}\alpha^{2}\left\lVert\nabla F(x_{k})\right\rVert}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}\leqslant L_{1}\alpha^{2}\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right].$
Then, we have the upper bound
$\quad\;-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\frac{L_{1}\alpha^{2}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}$ (C.13)
$\leqslant(\alpha+L_{1}\alpha^{2})\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right]$
$\leqslant 2\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right],$ (C.14)
since $\alpha\leqslant 1/L_{1}$.
By the A.M.-G.M. inequality (i.e., $2ab\leqslant\frac{1}{\rho_{1}}a^{2}+\rho_{1}b^{2}$ for any $\rho_{1}>0$), we have
 $\quad\;2\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right]$
$\leqslant\frac{\alpha}{\rho_{1}}\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\rho_{1}\alpha}{\sqrt{v_{k-1}}}\left(\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right]\right)^{\negthickspace 2}$
$\leqslant\frac{\alpha}{\rho_{1}}\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\rho_{1}\alpha}{\sqrt{v_{k-1}}}\mathbb{E}_{k}[\left\lVert g_{k}\right\rVert^{2}]\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{(\sqrt{v_{k}}+\sqrt{v_{k-1}})^{2}}\right]$ by Hölder’s inequality
$\leqslant\frac{\alpha}{\rho_{1}}\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\rho_{1}\alpha\tau}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{(\sqrt{v_{k}}+\sqrt{v_{k-1}})^{2}}\right].$
Also recall from the proof of Theorem 5.1 that
$\quad\;\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{(\sqrt{v_{k}}+\sqrt{v_{k-1}})^{2}}\right]$
$\leqslant\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]$
$=\mathbb{E}_{k}\left[\frac{\left\lVert\nabla F(x_{k-1})\right\rVert^{2}}{\sqrt{v_{k-1}}}-\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k}}}\right]+\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}-\left\lVert\nabla F(x_{k-1})\right\rVert^{2}}{\sqrt{v_{k-1}}}.$
Thus, by the reverse triangle inequality, we have
$\left\lVert\nabla F(x_{k})\right\rVert-\left\lVert\nabla F(x_{k-1})\right\rVert\leqslant\left\lVert\nabla F(x_{k})-\nabla F(x_{k-1})\right\rVert\leqslant(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert x_{k}-x_{k-1}\right\rVert.$
On the other hand, by the triangle inequality and Assumption 2, we have
$\left\lVert\nabla F(x_{k-1})\right\rVert$ $\leqslant\left\lVert\nabla F(x_{k-1})-\nabla F(x_{k})\right\rVert+\left\lVert\nabla F(x_{k})\right\rVert$
$\leqslant(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert x_{k}-x_{k-1}\right\rVert+\left\lVert\nabla F(x_{k})\right\rVert.$
Therefore, we obtain
$\quad\;\left\lVert\nabla F(x_{k})\right\rVert^{2}-\left\lVert\nabla F(x_{k-1})\right\rVert^{2}$
$=\left(\left\lVert\nabla F(x_{k})\right\rVert-\left\lVert\nabla F(x_{k-1})\right\rVert\right)\left(\left\lVert\nabla F(x_{k})\right\rVert+\left\lVert\nabla F(x_{k-1})\right\rVert\right)$
$\leqslant(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert x_{k}-x_{k-1}\right\rVert((L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert x_{k}-x_{k-1}\right\rVert+2\left\lVert\nabla F(x_{k})\right\rVert)$
$=(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+2(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert,$
implying that
$\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]\leqslant\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]\\ +\frac{1}{\sqrt{v_{k-1}}}\left[(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+2(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert\right].$
By the A.M.-G.M. inequality, we have $(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)^{2}\leqslant 2(L_{0}^{2}+L_{1}^{2}\left\lVert\nabla F(x_{k})\right\rVert^{2})$, which leads to
$\quad\;\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]$
$\leqslant\frac{1}{\sqrt{v_{k-1}}}\left[2(L_{0}^{2}+L_{1}^{2}\left\lVert\nabla F(x_{k})\right\rVert^{2})\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+2(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert\right]$
$\qquad\qquad+\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]$
$=\frac{2L_{0}}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert+\frac{2L_{1}}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}\left\lVert x_{k}-x_{k-1}\right\rVert$
$\qquad\qquad+\frac{2L_{0}^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\frac{2L_{1}^{2}}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}].$
Again, by the A.M.-G.M. inequality, with any $\rho_{2}>0$,
$\rho_{1}\alpha\tau\cdot\frac{2L_{0}}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert\leqslant\frac{\rho_{1}\alpha}{\rho_{2}\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}+\frac{\rho_{1}\rho_{2}L_{0}^{2}\tau^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}.$
The choice of $\alpha\leqslant\omega/(2L_{1}\rho_{1}\tau)$ and $\left\lVert x_{k}-x_{k-1}\right\rVert\leqslant\alpha$ implies
$\rho_{1}\alpha\tau\cdot\frac{2L_{1}}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}\left\lVert x_{k}-x_{k-1}\right\rVert\leqslant\frac{2L_{1}\rho_{1}\alpha^{2}\tau}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}\leqslant\omega\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}.$
Also, the choice of $\alpha\leqslant\sqrt{\omega}/(L_{1}\sqrt{2\rho_{1}\tau})$ and $\left\lVert x_{k}-x_{k-1}\right\rVert\leqslant\alpha$ implies
$\rho_{1}\alpha\tau\cdot\frac{2L_{1}^{2}}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}\leqslant\frac{2L_{1}^{2}\rho_{1}\alpha^{3}\tau}{\sqrt{v_{k-1}}}\left\lVert\nabla F(x_{k})\right\rVert^{2}\leqslant\omega\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}.$
Consequently, we have
$\rho_{1}\alpha\tau\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]\leqslant\alpha\left(\frac{\rho_{1}}{\rho_{2}}+2\omega\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{2L_{0}^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}\\ +\frac{\rho_{1}\rho_{2}L_{0}^{2}\tau^{2}}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\rho_{1}\alpha\tau\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}].$
Then, (C.14) is further upper bounded by
$-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\frac{L_{1}\alpha^{2}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}\\ \leqslant\alpha\left(\frac{1}{\rho_{1}}+\frac{\rho_{1}}{\rho_{2}}+2\omega\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{L_{0}^{2}\rho_{1}\tau(2\alpha+\rho_{2}\tau)}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\rho_{1}\alpha\tau\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}].$
Finally, (C.11) becomes
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})+\alpha\left(\frac{1}{\rho_{1}}+\frac{\rho_{1}}{\rho_{2}}+2\omega-1\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{L_{0}\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}\\ +\frac{L_{0}^{2}\rho_{1}\tau(2\alpha+\rho_{2}\tau)}{\sqrt{v_{k-1}}}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+\rho_{1}\alpha\tau\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}].$
Note that our choice of $(\rho_{1},\rho_{2},\omega)\in\mathbb{R}_{++}^{3}$ implies that $11/\rho_{1}+\rho_{1}/\rho_{2}+2\omega<$ 1 and $\alpha\leqslant 1/L_{1}$. Following similar arguments in the proof of Theorem 5.1, the result follows. ∎
Appendix D Details and Additional Results of Numerical Experiments
For the implementation of the proposed schemes, per-sample gradients are computed using JAX-like composable function transforms called torch.func in PyTorch 2.0+. Numerical experiments are carried out on workstations with NVIDIA RTX 2080Ti 11GB (for MNIST), A100 80GB (for CNN on CIFAR-10) and L40S 48GB (for ResNet-18 on CIFAR-10) GPUs, with PyTorch 2.2.1 (Paszke et al., 2019) and Lightning Fabric 2.2.0 (Falcon and The PyTorch Lightning team, 2019). The AdaGrad-Norm implementation is taken from (Ward et al., 2019).
D.1 Multi-class Logistic Regression on MNIST
We initially applied our methods to a ten-class logistic regression problem using the MNIST dataset, with a smooth convex objective. Our experiments were conducted with an equal training budget of 6 million samples (equivalent to 100 epochs), setting a maximum batch size of 60,000 (i.e., the full batch) for all approaches. To highlight the adaptivity and flexibility of our proposed methods, we refrained from conducting an exhaustive search for optimal values of $\alpha$, $\eta$, and $\vartheta$, and we did not employ any learning rate schedules across the methods. The outcomes, including the number of iterations required (steps), average batch sizes (bsz.), final training loss (loss), and final validation accuracy (acc.), are documented in Table 3.
Table 3: Multi-class logistic regression on MNIST
Scheme test steps bsz. loss acc.
AdaSGD $\eta=0.10$ 351 17131 1.04 0.82
AdaSGD $\eta=0.25$ 1029 5831 0.67 0.86
AdAdaGrad-Norm $\eta=0.10$ 596 10060 1.50 0.78
AdAdaGrad-Norm $\eta=0.25$ 2462 2437 1.02 0.83
AdAdaGrad $\eta=0.10$ 126 47282 0.54 0.88
AdAdaGrad $\eta=0.25$ 274 21918 0.46 0.90
AdaSGD $\vartheta=0.05$ 717 8362 0.76 0.85
AdaSGD $\vartheta=0.10$ 1804 3327 0.54 0.88
AdAdaGrad-Norm $\vartheta=0.05$ 1127 5322 1.28 0.80
AdAdaGrad-Norm $\vartheta=0.10$ 5349 1122 0.80 0.85
The results, coupled with the graphical representations in Figure 4, reveal that AdAdaGrad, utilizing the norm test with $\eta=0.25$, outperforms others in terms of both training loss and validation accuracy. These findings underscore the importance of adaptive (and coordinate-wise) learning rates within adaptive sampling methods. Despite the best performance of AdAdaGrad using the norm test with $\eta=0.25$, we note that the one with $\eta=0.10$ is more hardware-efficient—having an average batch size of more than 47,000, totaling only 126 steps (i.e., gradient evaluation) to run through all 6 million samples. This implies an intrinsic trade-off between computationally efficiency and generalization. Interestingly, AdAdaGrad-Norm, while theoretically simpler to analyze, might underperform compared to AdaSGD (i.e., SGD with adaptive batch size schemes) for this specific convex problem and set of hyperparameters. It is also noteworthy that the norm test tends to increase batch sizes more aggressively than the inner product test, leading to more efficient utilization of available GPU memory.
Figure 4: Training loss, validation accuracy and batch size curves (vs. number of training samples) of AdaSGD, AdAdaGrad and AdAdaGrad-Norm for logistic regression on the MNIST dataset.
Table 4: Training hyperparameters for multi-class logistic regression on MNIST
Model Multi-class Logistic Regression
Training samples 6M (100 epochs)
Weight initialization Default
Learning rate schedule None
Optimizer SGD or AdaGrad(-Norm)
Base learning rate 0.008
Base batch size 2
Maximum global batch size 60000
Weight decay 0
Momentum 0
Precision tf32
D.2 Three-layer Convolutional Neural Network on MNIST
Table 5: Three-layer CNN on MNIST
Scheme test steps time (h) bsz. loss acc.
SGD N/A 2929 6.77 2048 0.12 0.96
SGD N/A 1464 3.38 4096 0.20 0.94
SGD N/A 732 1.72 8192 0.32 0.91
SGD N/A 366 0.90 16384 0.51 0.87
SGD N/A 183 0.45 32768 1.54 0.75
SGD N/A 99 0.27 60000 2.15 0.66
AdaGrad N/A 2929 7.12 2048 0.02 0.99
AdaGrad N/A 1464 3.60 4096 0.02 0.99
AdaGrad N/A 732 1.82 8192 0.05 0.98
AdaGrad N/A 366 0.92 16384 0.07 0.98
AdaGrad N/A 199 0.52 30000 0.10 0.97
AdaGrad N/A 183 0.47 32768 0.11 0.97
AdaGrad N/A 149 0.36 40000 0.13 0.96
AdaGrad N/A 99 0.29 60000 0.17 0.95
AdaSGD norm; $\eta=0.10$ 256 0.73 23546 0.79 0.83
AdaSGD norm; $\eta=0.25$ 383 1.05 15627 0.48 0.88
AdAdaGrad-Norm norm; $\eta=0.10$ 226 0.65 26567 0.88 0.83
AdAdaGrad-Norm norm; $\eta=0.25$ 435 1.27 13830 0.54 0.87
AdAdaGrad norm; $\eta=0.10$ 149 0.45 40057 0.15 0.96
AdAdaGrad norm; $\eta=0.25$ 198 0.58 30152 0.13 0.97
AdAdaGrad norm; $\eta=0.5$ 215 0.62 27940 0.11 0.97
AdAdaGrad norm; $\eta=0.75$ 271 0.79 22228 0.10 0.97
AdaSGD inner product; $\vartheta=0.01$ 230 0.63 26078 0.98 0.80
AdaSGD inner product; $\vartheta=0.05$ 411 1.17 14593 0.45 0.88
AdAdaGrad-Norm inner product; $\vartheta=0.01$ 241 0.70 24872 0.83 0.84
AdAdaGrad-Norm inner product; $\vartheta=0.05$ 528 1.44 11365 0.50 0.88
Table 6: Training hyperparameters for three-layer CNN on MNIST
Model 3-layer CNN on MNIST
Training samples 6M (100 epochs)
Weight initialization Default
Training samples 6M (100 epochs)
Optimizer SGD or AdaGrad(-Norm)
Base learning rate 0.008
Base batch size 8
Maximum batch size 60000
Weight decay 0
Momentum 0
Precision tf32
Figure 5: Training loss, validation accuracy and batch size curves (vs. number of training samples) of AdaSGD, AdAdaGrad and AdAdaGrad-Norm for three-layer CNN on the MNIST dataset.
Figure 6: Training loss, validation accuracy and batch size curves (vs. number of training samples) of SGD and AdaSGD for three-layer CNN on the MNIST dataset.
Figure 7: Training loss, validation accuracy and batch size curves (vs. number of training samples) of AdaGrad and AdAdaGrad for three-layer CNN on the MNIST dataset.
D.3 Three-layer Convolutional Neural Network on CIFAR-10
We then consider a similar problem with a three-layer CNN with a slightly different structure on the harder CIFAR-10 dataset. We use a training budget of 5 million samples (100 epochs) and a maximum batch size of 10,000 samples. To reduce the computational overhead brought by the tests, we perform the test every 10 steps. In Figure 8, we again observe that AdaSGD converges more slowly than AdAdaGrad and AdAdaGrad-Norm. This would be due to the lack of a well-crafted learning rate scaling with respect to batch sizes (cf. scaling rule) for SGD in the nonconvex case: the rapid increase in its batch sizes implies a very small effective learning rate, which is equal to the ratio of the learning rate and the batch size. Without proper rescaling of the learning rate, such a small effective learning rate would potentially decelerate convergence. We could empirically confirm this claim since AdaSGD using the inner product test with $\vartheta=0.1$ has its batch size increasing slowly and eventually plateaus at somewhere below 4,000. It also achieves much faster convergence compared to its SGD counterparts, with a final validation accuracy of 57%, approaching the performance of adaptive methods. We however point out that this AdaSGD instance has a much smaller average batch size, hence taking much more gradient updates than the adaptive methods given an equal budget of training samples.
Table 7: Training hyperparameters for three-layer CNN on CIFAR-10
Model 3-layer CNN on CIFAR-10
Training samples 5M (100 epochs)
Weight initialization Default
Learning rate schedule None
Training samples 5M (100 epochs)
Optimizer SGD or AdaGrad(-Norm)
Optimizer scaling rule None
Base learning rate 0.05
Base batch size 2
Maximum batch size 10000
Weight decay 0
Momentum 0
Precision tf32
Table 8: Three-layer CNN on CIFAR-10
Scheme test steps bsz. loss acc.
AdaSGD norm; $\eta=0.25$ 523 9544 1.68 0.40
AdaSGD norm; $\eta=0.50$ 658 7592 1.59 0.43
AdAdaGrad-Norm norm; $\eta=0.25$ 531 9401 1.36 0.52
AdAdaGrad-Norm norm; $\eta=0.50$ 1261 3964 1.19 0.57
AdAdaGrad norm; $\eta=0.25$ 903 5533 1.20 0.54
AdAdaGrad norm; $\eta=0.50$ 1123 4451 1.11 0.57
AdaSGD inner product; $\vartheta=0.05$ 1597 3130 1.17 0.57
AdaSGD inner product; $\vartheta=0.10$ 640 7806 1.64 0.41
AdAdaGrad-Norm inner product; $\vartheta=0.05$ 780 6413 1.29 0.55
AdAdaGrad-Norm inner product; $\vartheta=0.10$ 1948 2567 1.13 0.58
Figure 8: Training loss, validation accuracy and batch size curves (vs. number of training samples) of AdaSGD, AdAdaGrad and AdAdaGrad-Norm for three-layer CNN on the CIFAR-10 dataset.
D.4 ResNet-18 on CIFAR-10
Table 9: Training hyperparameters for ResNet-18 on CIFAR-10
Model ResNet-18 on CIFAR-10
Training samples 10M (200 epochs)
Weight initialization Default
Optimizer AdaGrad or Adam
Learning rate schedule Linear warmup + cosine decay
Learning rate warmup (samples) 1M
$(\beta_{1},\beta_{2})$ $(0.9,0.95)$
$\varepsilon$ $10^{-8}$
Peak learning rate 0.05
Minimum learning rate 0.005
Base batch size 8
Maximum batch size 50000
Weight decay 0
Precision tf32
Table 10: ResNet-18 on CIFAR-10
Scheme test steps time (h) bsz. loss acc.
AdaGrad N/A 2441 0.88 4096 0.0042 0.8521
AdaGrad N/A 1220 0.70 8192 0.0808 0.8072
AdaGrad N/A 610 0.56 16384 0.5098 0.7264
AdaGrad N/A 305 0.32 32768 0.9684 0.5816
AdaGrad N/A 199 0.23 50000 1.3625 0.4708
Adam N/A 2441 1.20 4096 0.0003 0.9147
Adam N/A 1220 0.97 8192 0.0004 0.8946
Adam N/A 610 0.77 16384 0.0028 0.8628
Adam N/A 305 0.45 32768 0.4000 0.7463
Adam N/A 199 0.33 50000 1.0680 0.5750
AdAdaGrad norm; $\eta=0.025$ 222 0.32 44934 1.2770 0.5107
AdAdaGrad norm; $\eta=0.05$ 485 0.60 20615 0.6204 0.7079
AdAdaGrad norm; $\eta=0.075$ 1697 1.02 5892 0.0258 0.8180
AdAdaGrad norm; $\eta=0.1$ 1404 0.94 7123 0.0668 0.8085
AdAdam norm; $\eta=0.025$ 211 0.34 47380 0.9039 0.6234
AdAdam norm; $\eta=0.05$ 426 0.60 23463 0.0061 0.8228
AdAdam norm; $\eta=0.075$ 900 0.74 11108 0.0008 0.8983
AdAdam norm; $\eta=0.1$ 1126 0.82 8880 0.0000 0.9042